OpenStudy (anonymous):

Mark any solutions to the equation 2cos2x - 1 = 0.

hero (hero):

Did you mean \(2 \cos(2x) - 1 = 0\) or \(2 \cos^2(x) - 1 = 0\)

OpenStudy (anonymous):

2cos^2(x)-1=0

OpenStudy (marissalovescats):

Well you solve the equation for cos (x) :